Write the equation of the quadratic function with roots -9 and and -3 and a vertex at (-6, -1)
Answer by Theo(13342)

if the roots are -9 and -3, then the factors are (x+9) * (x+3) = 0
multiply those factors together to get:
x^2 + 12x + 27 = 0

the standard form of the quadratic is ax^2 + b + c = 0
the formula for the vertex of a quadratic equation is x = -b/2a.

in the equation of f(x) = x^2 + 12x + 27
a = 1
b = 12
c = 27

using those values, you would get the x-coordinate of the vertex as:
x = -b/2a = -12/2 = -6.

when x = -6, the value of y = (-6)^2 + 12(-6) + 27 = 36 - 72 + 27 = 63 - 72 = -9.

the value of y is equal to -9.
you need it to be -1.

if you multiply -9 by 1/9, it will be come -1.

so you multiply both sides of the equation of x^2 + 12x + 27 = 0 by 1/9 and you get:

1/9 * (x^2 + 12x + 27) = 1/9 * 0

this results in:

(1/9)*x^2 + (12/9)*x + 27/9 = 0

simplify this to get:

(1/9)x^2 + (4/3)x + 3 = 0

that should be your equation.

now solve for x = -b/2a to get the x-coordinate of the vertex.

x = -b/2a becomes x = (-4/3) / (2/9) which becomes x = (-4/3) * (9/2) which becomes x = -36 / 6 which becomes x = -6.

the x-coordinate of the vertex is still at -6.

now solve for the y-coordinate of the vertex.

f(x) = (1/9)x^2 + (4/3)x + 3 is evaluated at x = -6.

you get f(-6) = (1/9)(36) + (4/3)(-6) + 3 which becomes:

f(-6) = 4 - 8 + 3 which becomes:

f(-6) = -1

now you have the equation where you want it.

you can multiply any equation by a factor and the equation will stretch or shrink along the y-axis without changing the roots, depending on the fraction.

in this case, the original equation of x^2 + 12x + 27 = 0 was shrunk because you multiplied both sides of the equation by 1/9.

the attached graphs will show you what i mean.

here's the equation of x^2 + 12x + 27 as it original was without any multiplication.
when y = 0, the value of x is equal to -3 and -9.

$$$

here's the equation of (1/9)x^2 + (4/3)x + 3.
this is the original equation multiplied by (1/9).
the result is that the equation shrinks along the y-axis.
it becomes shorter in both directions from the x-axis.\
the vertex that was at -6,-9 is not as -6,-1.
the roots remain the same.

$$$

here's both equations shown together so you can see how the revised graph looks in relationship to the orgiinal graph.
the original graph has a vertex at (-6,-9).
the revised graph has a vertex at (-6,-1).

$$$